# On the non-Markovian quantum control dynamics

### v2

@@ II-A Single multi-level atom coupled with the cavity
$A_{n}(t)=\begin{bmatrix}-\left(F(t)+F^{*}(t)\right)|\kappa_{n}|^{2}&-ig_{n}e^{i\delta t}&ig_{n}e^{-i\delta t}\\ -ig_{n}e^{-i\delta t}&-i\delta-F^{*}(t)|\kappa_{n}|^{2}&0\\ ig_{n}e^{i\delta t}&0&i\delta-F(t)|\kappa_{n}|^{2}\end{bmatrix},$ (13)
$B_{n}(t)=\begin{bmatrix}0&ig_{n+1}e^{i\delta t}&-ig_{n+1}e^{-i\delta t}\\ 0&0&0\\ 0&0&0\end{bmatrix},$ (14)
$C_{n}(t)$ $=\begin{bmatrix}0&ig_{n}e^{i\delta t}&-ig_{n}e^{-i\delta t}\end{bmatrix},$ (15)
$D_{n}(t)$ $=\begin{bmatrix}0\\ ig_{n}e^{-i\delta t}\\ -ig_{n}e^{i\delta t}\end{bmatrix},$ (16)
@@
The analysis in Section II and Section III is based on the following assumption for initial settings.
Assumption 3.
Assume initially the atom is excited at the highest energy level and $\langle\sigma^{+}_{n}a\rangle=\langle\sigma^{-}_{n}a^{{\dagger}}\rangle=\langle a^{{\dagger}}a\rangle=0$.
Remark 2.
Eq. (12) is a time-varying linear equation and reduces to the time-invariant case only when $\delta=0$ and $F(t)$ becomes a constant, thereby converging to the Markovian case.
In the following subsection, we clarify the dynamics of $F(t)$ from the perspective of nonlinear dynamics, and then analyze how the time-varying $F(t)$ can influence the quantum control performance.
@@ II-B Non-Markovian parameter dynamics
$P=\sum_{n}|\kappa_{n}|^{2},$ (17a)
$Q=-\left(\gamma+i\Omega-i\omega_{a}\right),$ (17b)
$R=\frac{\gamma\chi}{2},$ (17c)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}F(t)=PF^{2}(t)+QF(t)+R,$ (18)
and $F(t)$ can be solved according to following different parameter settings.
II-B1 $\Omega=\omega_{a}$ and $\frac{R}{P}-\left(\frac{Q}{2P}\right)^{2}<0$
When $\Omega=\omega_{a}$ and $\frac{R}{P}-\left(\frac{Q}{2P}\right)^{2}<0$, namely $4PR<Q^{2}$ and $2\chi\sum_{n}|\kappa_{n}|^{2}<\gamma$, which means that the coupling between the atom and the environment can be relatively small and bounded by $\gamma$, then
$F(t)=\frac{2\sqrt{\left(\frac{Q}{2P}\right)^{2}-\frac{R}{P}}}{1-e^{C+\sqrt{Q^{2}-4PR}t}}-\left(\frac{Q}{2P}+\sqrt{\left(\frac{Q}{2P}\right)^{2}-\frac{R}{P}}\right),$ (19)
where $C$ is a constant and can be determined as
$e^{C}=1-2\sqrt{\left(\frac{Q}{2P}\right)^{2}-\frac{R}{P}}\left[\frac{Q}{2P}+\sqrt{\left(\frac{Q}{2P}\right)^{2}-\frac{R}{P}}\right]^{-1},$ (20)
based on the initial condition that $F(0)=0$.
II-B2 $\Omega=\omega_{a}$ and $\frac{R}{P}-\left(\frac{Q}{2P}\right)^{2}>0$
When $\Omega=\omega_{a}$ and $\frac{R}{P}-\left(\frac{Q}{2P}\right)^{2}>0$, namely $2\chi\sum_{n}|\kappa_{n}|^{2}>\gamma$, which means that the coupling between the atom and environment is relatively stronger. Then by Eq. (18),
$F(t)=\sqrt{\frac{R}{P}-\left(\frac{Q}{2P}\right)^{2}}\tan\left[\sqrt{\frac{R}{P}-\left(\frac{Q}{2P}\right)^{2}}\left(Pt+C\right)\right]-\frac{Q}{2P},$ (21)
where $C$ can be similarly determined by $F(0)=0$ and the solution is not unique.
II-B3 $\Omega=\omega_{a}$ and $\frac{R}{P}-\left(\frac{Q}{2P}\right)^{2}=0$
When $\Omega=\omega_{a}$ and $\frac{R}{P}-\left(\frac{Q}{2P}\right)^{2}=0$, namely $2\chi\sum_{n}|\kappa_{n}|^{2}=\gamma$. Then by Eq. (18), $C=-2P/Q$ and
$F(t)=-\frac{Q}{2P}-\frac{1}{Pt+C}.$ (22)
We have the following proposition for the relationship between the parameter settings above and the non-Markovian dynamics.
Proposition 1.
When $\Omega=\omega_{a}$ and $\frac{R}{P}-\left(\frac{Q}{2P}\right)^{2}\leq 0$, the master equation for open quantum system finally converges to the Markovian format when $t\rightarrow\infty$.
@@ II-B3 $\Omega=\omega_{a}$ and $\frac{R}{P}-\left(\frac{Q}{2P}\right)^{2}=0$
According to the calculations above, when $\Omega=\omega_{a}$ and $\frac{R}{P}-\left(\frac{Q}{2P}\right)^{2}\leq 0$,
$\lim_{t\rightarrow\infty}F(t)=-\left(\frac{Q}{2P}+\sqrt{\left(\frac{Q}{2P}\right)^{2}-\frac{R}{P}}\right)\triangleq F_{\infty}.$ (23)
When $\frac{R}{P}-\left(\frac{Q}{2P}\right)^{2}>0$, $\lim_{t\rightarrow\infty}F(t)$ doesn’t exist because $\tan\left[\sqrt{\frac{R}{P}-\left(\frac{Q}{2P}\right)^{2}}\left(Pt+C\right)\right]$ is infinite when
$\sqrt{\frac{R}{P}-\left(\frac{Q}{2P}\right)^{2}}\left(Pt+C\right)=n\pi+\pi/2$
with $n=0,1,2,\cdots$. ∎
For the general case that $F(t)$ is complex when $\Omega\neq\omega_{a}$, we denote $F(t)=F_{R}(t)+iF_{I}(t)$ with $F_{R}(t)$ and $F_{I}(t)$ representing the real and imaginary part of $F(t)$ respectively, then we can derive the following real-valued nonlinear equation
$\frac{\mathrm{d}}{\mathrm{d}t}F_{R}(t)=\sum_{n}|\kappa_{n}|^{2}\left(F_{R}^{2}(t)-F_{I}^{2}(t)\right)-\gamma F_{R}(t)$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\left(\Omega-\omega_{a}\right)F_{I}(t)+\frac{\gamma\chi}{2},$ (24a)
$\frac{\mathrm{d}}{\mathrm{d}t}F_{I}(t)=2\sum_{n}|\kappa_{n}|^{2}F_{R}(t)F_{I}(t)-\left(\Omega-\omega_{a}\right)F_{R}(t)-\gamma F_{I}(t).$ (24b)
Denote $X_{F}(t)=\left[F_{R}(t),F_{I}(t)\right]^{T}$ and $X_{F}(0)=\left[0,0\right]^{T}$. Then Eq. (24) can be rewritten as
$\frac{\mathrm{d}}{\mathrm{d}t}X_{F}(t)$ $=\begin{bmatrix}\sum_{n}|\kappa_{n}|^{2}F_{R}(t)-\gamma&-\sum_{n}|\kappa_{n}|^{2}F_{I}(t)\\ \sum_{n}|\kappa_{n}|^{2}F_{I}(t)&\sum_{n}|\kappa_{n}|^{2}F_{R}(t)-\gamma\end{bmatrix}\begin{bmatrix}F_{R}(t)\\ F_{I}(t)\end{bmatrix}$ (25)
$~{}~{}~{}~{}+\begin{bmatrix}0&\left(\Omega-\omega_{a}\right)\\ -\left(\Omega-\omega_{a}\right)&0\end{bmatrix}\begin{bmatrix}F_{R}(t)\\ F_{I}(t)\end{bmatrix}+\begin{bmatrix}\gamma\chi/2\\ 0\end{bmatrix}$
@@
where $\textbf{f}\left(X_{F}(t)\right)$ represents the sum of the first and third terms after the first equal sign, $\textbf{f}_{\Omega}=\begin{bmatrix}0&\left(\Omega-\omega_{a}\right)\\ -\left(\Omega-\omega_{a}\right)&0\end{bmatrix}$.
Based on the contraction analysis for arbitrary initial condition of $X_{F}(t)$ [62], consider the differential relation in Eq. (25), then
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\left(\delta X_{F}(t)^{T}\delta X_{F}(t)\right)$ $=2\delta X_{F}(t)^{T}\delta\dot{X}_{F}(t)$ (27)
$=2\delta X_{F}(t)^{T}\left[\frac{\partial\textbf{f}}{\partial X_{F}}\left(X_{F}(t)\right)+\textbf{f}_{\Omega}\right]\delta X_{F}(t)$
$=2\delta X_{F}(t)^{T}\frac{\partial\textbf{f}}{\partial X_{F}}\delta X_{F}(t)+2\textbf{f}_{\Omega}\delta X_{F}(t)^{T}\delta X_{F}(t).$
The stability of Eq. (25) is mainly determined by the Jacobian $\textbf{J}=\frac{\partial\textbf{f}}{\partial X_{F}}+\textbf{f}_{\Omega}$, according to Eq. (25),
$\textbf{J}\left(X_{F}(t)\right)=\begin{bmatrix}PF_{R}(t)-\gamma&-PF_{I}(t)+\left(\Omega-\omega_{a}\right)\\ PF_{I}(t)-\left(\Omega-\omega_{a}\right)&PF_{R}(t)-\gamma\end{bmatrix},$ (28)
which reduces to $\frac{\partial\textbf{f}}{\partial X_{F}}$ when $\Omega=\omega_{a}$.
For the nonlinear system in Eq. (24), we regard $u=\Omega-\omega_{a}$ as the control,
@@
and define the output of the system as
$y_{F}(t)$ $=\begin{bmatrix}1&1\\ 1&-1\end{bmatrix}\begin{bmatrix}F_{R}(t)\\ F_{I}(t)\end{bmatrix}\triangleq CX_{F}(t).$ (30)
Because the norm of matrix $C$ is finite, the norm of the output vector $y_{F}(t)$ is determined by the state vector $X_{F}(t)$. We further analyze the control input and output properties according to the following definition.
Definition 2.
(BIBO) [63] The system in Eq. (29) is said to be bounded-input bounded-output (BIBO) if for every $a_{u}\geq 0$ and $\alpha_{X}\geq 0$ there is a finite number $\beta_{X}=\beta_{X}\left(a_{u},\alpha_{X}\right)$ such that for every initial condition $X_{F}\left(t_{0}\right)$ with $X_{F}\left(t_{0}\right)\leq\alpha_{X}$ and every control sequence $u$ with $\|u\|\leq a_{u}$,
for $t^{\prime}>t_{0}$.
Additionally, a discrete generalization of Definition 2 is given in Ref. [64].
@@
$\frac{\mathrm{d}}{\mathrm{d}t}X_{F}(t)$ $=\mathbf{A}\left(X_{F}(t)\right)X_{F}(t)+\mathbf{B}X_{F}(t)u+\begin{bmatrix}\gamma\chi/2\\ 0\end{bmatrix}$ (31)
$=\left(\mathbf{A}\left(X_{F}(t)\right)+\mathbf{B}u\right)X_{F}(t)+\begin{bmatrix}\gamma\chi/2\\ 0\end{bmatrix},$
where $\textbf{f}\left(X_{F}(t)\right)=\mathbf{A}\left(X_{F}(t)\right)X_{F}(t)+\begin{bmatrix}\gamma\chi/2\\ 0\end{bmatrix}$, $\mathbf{B}=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$. In the second line of Eq. (31), the parameter $u$ can be regarded as a constant perturbation or random unknown uncertainty when the environment cannot be precisely modeled.
According to Refs. [65, 62], the Lyapunov function can be defined as $V\left(X_{F}\right)=\textbf{f}^{T}\left(X_{F}\right)\mathbf{M}\left(X_{F}\right)\textbf{f}\left(X_{F}\right)$, where $\mathbf{M}\left(X_{F}\right)$ is a contraction matrix to be determined. When $u=0$,
$\dot{V}(X_{F})=\textbf{f}^{T}\left(X_{F}(t)\right)\left[\frac{\partial\textbf{f}^{T}}{\partial X_{F}}\mathbf{M}\left(X_{F}\right)+\mathbf{M}\left(X_{F}\right)\frac{\partial\textbf{f}}{\partial X_{F}}+\dot{\mathbf{M}}\right]\textbf{f}\left(X_{F}(t)\right).$ (32)
For the simplest case with $\Omega=\omega_{a}$ and $F_{I}(t)\equiv 0$, Eq. (25) reduces to $\frac{\mathrm{d}}{\mathrm{d}t}F_{R}(t)=PF_{R}^{2}(t)-\gamma F_{R}(t)+\frac{\gamma\chi}{2}.$ More explanations are given combined with the following example and propositions.

### v3

@@ II-A Single multi-level atom coupled with the cavity
$A_{n}=$ $-\kappa_{n}{\rm{diag}}\left[(F_{n}+F_{n}^{*}),F_{n}^{*},F_{n}\right]$ (13)
$+\begin{bmatrix}0&-i\tilde{g}_{n}&i\tilde{g}_{n}^{*}\\ -i\tilde{g}_{n}^{*}&-i\Delta_{n}&0\\ i\tilde{g}_{n}&0&i\Delta_{n}\end{bmatrix},$
$B_{n}=\begin{bmatrix}\kappa_{n+1}\left(F_{n+1}+F_{n+1}^{*}\right)&i\tilde{g}_{n+1}&-i\tilde{g}_{n+1}^{*}\\ 0&0&0\\ 0&0&0\end{bmatrix},$ (14)
$C_{n}$ $=\begin{bmatrix}0&i\tilde{g}_{n}&-i\tilde{g}_{n}^{*}\end{bmatrix},$ (15)
$D_{n}$ $=\begin{bmatrix}0&i\tilde{g}_{n}^{*}&-i\tilde{g}_{n}\end{bmatrix}^{\rm T},$ (16)
@@
For the basic control model of this paper in Eq. (12), which is a time-varying linear equation, the state vector $X(t)$ in Eq. (11) is made up of the atomic state and photonic state in the cavity. The target of this paper is to study the condition for the occurrence of non-Markovian interactions between quantum system and environment evaluated by the nonlinear dynamics in Eq. (9e), and how the stability of time-varying atomic and photonic dynamics stability in Eq. (12) can be influenced by the non-Markovian process and additional control methods in the following.
@@ II-B Non-Markovian parameter dynamics
In this subsection, we clarify the dynamics of $F_{n}(t)$ from the perspective of nonlinear dynamics, and then analyze how the time-varying $F_{n}(t)$ can influence the quantum control performance in Eq. (9) in the following Sec. III.
@@
$Q_{n}=-\left(\gamma+i\Omega-i\tilde{\omega}_{n}\right),$ (17a)
$S_{n}=\frac{\gamma\chi_{n}}{2},$ (17b)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}F_{n}(t)=\kappa_{n}F_{n}^{2}(t)+Q_{n}F_{n}(t)+S_{n},$ (18)
where $F_{n}(t)$ can be solved according to following different parameter settings.
II-B1 $\Omega=\tilde{\omega}_{n}$ and $04\kappa_{n}S_{n}-Q_{n}^{2}<$ 0
When the conditions are satisfied, namely $4\kappa_{n}S_{n}<Q_{n}^{2}$ and $2\chi_{n}\kappa_{n}<\gamma$, which means that the coupling between the atom and the environment can be relatively small and bounded by $\gamma$, then
$F_{n}=\frac{2\sqrt{\left(\frac{Q_{n}}{2\kappa_{n}}\right)^{2}-\frac{S_{n}}{\kappa_{n}}}}{1-e^{C+\sqrt{Q_{n}^{2}-4\kappa_{n}S_{n}}t}}-\frac{Q_{n}+\sqrt{Q_{n}^{2}-4\kappa_{n}S_{n}}}{2\kappa_{n}},$ (19)
where $C$ is a constant and can be determined by $F_{n}(0)$. For example, $e^{C}=1-2\sqrt{Q_{n}^{2}-4S_{n}\kappa_{n}}\left[Q_{n}+\sqrt{Q_{n}^{2}-4S_{n}\kappa_{n}}\right]^{-1}$ if $F_{n}(0)=0$.
II-B2 $\Omega=\tilde{\omega}_{n}$ and $04\kappa_{n}S_{n}-Q_{n}^{2}>$ 0
In this case, $2\chi_{n}\kappa_{n}>\gamma$, which means that the coupling between the atom and environment is relatively stronger. Then by Eq. (18),
$F_{n}=$ $-\frac{Q_{n}}{2\kappa_{n}}+\sqrt{\frac{S_{n}}{\kappa_{n}}-\left(\frac{Q_{n}}{2\kappa_{n}}\right)^{2}}$ (20)
×$\times\tan\left[\sqrt{\frac{S_{n}}{\kappa_{n}}-\left(\frac{Q_{n}}{2\kappa_{n}}\right)^{2}}\left(\kappa_{n}t+C\right)\right],$
where $C$ can be similarly determined by $F_{n}(0)$ and the solution is not unique.
II-B3 $\Omega=\tilde{\omega}_{n}$ and $04\kappa_{n}S_{n}-Q_{n}^{2}=$ 0
In this case, $2\chi_{n}\kappa_{n}=\gamma$, and
$F_{n}=-\frac{Q_{n}}{2\kappa_{n}}-\frac{1}{\kappa_{n}t+C},$ (21)
with $C=-2\kappa_{n}/Q_{n}$ if $F_{n}(0)=0$.
We have the following theorem for the relationship between the parameter settings above and the non-Markovian dynamics.
Theorem 1.
When $\tilde{\omega}_{n}=\Omega$ and $4\kappa_{n}S_{n}-Q_{n}^{2}\leq 0$, $F_{n}(t)$ in Eq. (18) converges to a constant when $t\rightarrow\infty$, resulting the trajectories for the population of atomic eigenstates related to the $n$-th level, namely $\langle\sigma^{+}_{n}\sigma^{-}_{n}\rangle$, $\langle\sigma^{+}_{n}a\rangle$ and $\langle\sigma^{-}_{n}a^{{\dagger}}\rangle$ in Eq. (9), finally converge to be Markovian when $t\rightarrow\infty$.
@@ II-B3 $\Omega=\tilde{\omega}_{n}$ and $04\kappa_{n}S_{n}-Q_{n}^{2}=$ 0
According to the calculations above, when $\Omega=\tilde{\omega}_{n}$ and $4\kappa_{n}S_{n}-Q_{n}^{2}\leq 0$,
$\lim_{t\rightarrow\infty}F_{n}(t)=-\frac{Q_{n}+\sqrt{Q_{n}^{2}-4\kappa_{n}S_{n}}}{2\kappa_{n}}.$ (22)
When $04\kappa_{n}S_{n}-Q_{n}^{2}>$ 0, $\lim_{t\rightarrow\infty}F_{n}(t)$ does not exist because $\tan\left[\sqrt{\frac{S_{n}}{\kappa_{n}}-\left(\frac{Q_{n}}{2\kappa_{n}}\right)^{2}}\left(\kappa_{n}t+C\right)\right]$ is infinite when
$\sqrt{\frac{S_{n}}{\kappa_{n}}-\left(\frac{Q_{n}}{2\kappa_{n}}\right)^{2}}\left(\kappa_{n}t+C\right)=l\pi+\pi/2,$
with $l=0,1,2,\cdots$. ∎
Based on this, we further generalize to the dynamics of arbitrary energy levels according to the stability of nonlinear dynamics in the following.
For the general case that $F_{n}(t)$ is complex when $\Omega\neq\tilde{\omega}_{n}$, we denote $F_{n}(t)=R_{n}(t)+iI_{n}(t)$ with $R_{n}(t)$ and $I_{n}(t)$ representing the real and imaginary part of $F_{n}(t)$ respectively, then we can derive the following real-valued nonlinear equations
$\frac{\mathrm{d}R_{n}}{\mathrm{d}t}=\kappa_{n}\left(R_{n}^{2}-I_{n}^{2}\right)-\gamma R_{n}+\left(\Omega-\tilde{\omega}_{n}\right)I_{n}+\frac{\gamma\chi_{n}}{2},$ (23a)
$\frac{\mathrm{d}I_{n}}{\mathrm{d}t}=2\kappa_{n}R_{n}I_{n}-\left(\Omega-\tilde{\omega}_{n}\right)R_{n}-\gamma I_{n}.$ (23b)
Denote $X_{F}(t)=\left[R_{n}(t),I_{n}(t)\right]^{\rm T}$ with $X_{F}(0)=\left[0,0\right]^{\rm T}$. Then Eq. (23) can be rewritten as
$\frac{\mathrm{d}}{\mathrm{d}t}X_{F}(t)=\begin{bmatrix}\kappa_{n}R_{n}-\gamma&-\kappa_{n}I_{n}\\ \kappa_{n}I_{n}&\kappa_{n}R_{n}-\gamma\end{bmatrix}\begin{bmatrix}R_{n}\\ I_{n}\end{bmatrix}$ (24)
$+\begin{bmatrix}0&\left(\Omega-\tilde{\omega}_{n}\right)\\ -\left(\Omega-\tilde{\omega}_{n}\right)&0\end{bmatrix}\begin{bmatrix}R_{n}(t)\\ I_{n}(t)\end{bmatrix}+\begin{bmatrix}\frac{\gamma\chi_{n}}{2}\\ 0\end{bmatrix}$
@@
where $\textbf{f}\left(X_{F}(t)\right)$ represents the sum of the first and third terms after the first equal sign, and $\textbf{f}_{\Omega}=\begin{bmatrix}0&\left(\Omega-\tilde{\omega}_{n}\right)\\ -\left(\Omega-\tilde{\omega}_{n}\right)&0\end{bmatrix}$.
Based on the contraction analysis for arbitrary initial condition of $X_{F}(t)$ [49], consider the differential in Eq. (24), then
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\left(\delta X_{F}(t)^{\rm T}\delta X_{F}(t)\right)$ (26)
$=$ $2\delta X_{F}(t)^{\rm T}\frac{\partial\textbf{f}}{\partial X_{F}}\delta X_{F}(t)+2\textbf{f}_{\Omega}\delta X_{F}(t)^{\rm T}\delta X_{F}(t).$
The stability of Eq. (24) is mainly determined by the Jacobian $\textbf{J}=\partial\textbf{f}/\partial X_{F}+\textbf{f}_{\Omega}$. According to Eq. (24),
$\textbf{J}\left(X_{F}\right)=\begin{bmatrix}\kappa_{n}R_{n}-\gamma&-\kappa_{n}I_{n}+\left(\Omega-\tilde{\omega}_{n}\right)\\ \kappa_{n}I_{n}-\left(\Omega-\tilde{\omega}_{n}\right)&\kappa_{n}R_{n}-\gamma\end{bmatrix},$ (27)
which reduces to $\partial\textbf{f}/\partial X_{F}$ when $\Omega=\tilde{\omega}_{n}$.
For the nonlinear system in Eq. (23), we regard $u=\Omega-\tilde{\omega}_{n}$ as an unknown input, which can be regarded as a constant perturbation or random unknown uncertainty due to the fact that the environment cannot be precisely modeled. Then
@@
and we define the output of the system as
$y_{F}(t)$ $=\begin{bmatrix}1&1\\ 1&-1\end{bmatrix}\begin{bmatrix}R_{n}(t)\\ I_{n}(t)\end{bmatrix}\triangleq\mathcal{C}X_{F}(t).$ (29)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}X_{F}(t)$ $=\left(\mathbf{A}\left(X_{F}(t)\right)+\mathbf{B}u\right)X_{F}(t)+\begin{bmatrix}\gamma\chi_{n}/2\\ 0\end{bmatrix},$ (30)
where $\textbf{f}\left(X_{F}(t)\right)=\mathbf{A}\left(X_{F}(t)\right)X_{F}(t)+\begin{bmatrix}\gamma\chi_{n}/2\\ 0\end{bmatrix}$, $\mathbf{B}=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$, and $\mathbf{A}$ is the first matrix on the RHS of Eq. (24).
According to Refs. [50, 49], the Lyapunov function can be defined as $V\left(X_{F}\right)=\textbf{f}^{\rm T}\left(X_{F}\right)\mathbf{M}\left(X_{F}\right)\textbf{f}\left(X_{F}\right)$, where $\mathbf{M}\left(X_{F}\right)$ is a contraction matrix to be determined. When $u=0$,
$\dot{V}(X_{F})=$ $\textbf{f}^{\rm T}\left(X_{F}(t)\right)\left[\frac{\partial\textbf{f}^{\rm T}}{\partial X_{F}}\mathbf{M}\left(X_{F}\right)\right.$ (31)
$\left.+\mathbf{M}\left(X_{F}\right)\frac{\partial\textbf{f}}{\partial X_{F}}+\dot{\mathbf{M}}\right]\textbf{f}\left(X_{F}(t)\right).$
For the simplest case with $\Omega=\tilde{\omega}_{n}$ and $I_{n}(t)\equiv 0$, Eq. (24) reduces to $\mathrm{d}R_{n}/\mathrm{d}t=\kappa_{n}R_{n}^{2}-\gamma R_{n}+\gamma\chi_{n}/2$. More explanations are given combined with the following example and propositions.